# New Perspectives on the Polyak Stepsize: Surrogate Functions and Negative Results

### v1

@@ Appendix D Additional Convergence Result for Generalized Polyak Stepsize
$C^{-1}(y)$ =$=\begin{cases}2y,&\text{ if }y\geq(\gamma K_{\nu})^{\frac{1+\nu}{1-\nu}}\\ (2\gamma K_{\nu}y)^{\frac{1+\nu}{2}},&\text{ if }y<(\gamma K_{\nu})^{\frac{1+\nu}{1-\nu}}\end{cases}$
@@ Appendix E Proofs for Section 6
In other words, the set of initializations that can converge to the optimal solution has measure zero.
@@
We divide $\mathbb{R}^{n}$ into two sets, $S=\bigcup_{k=1}^{\infty}T^{-k}(\{\boldsymbol{x}^{\star}\})$, and its compliment $S^{c}$. $S$ represents the points that can exactly reach the unique minimizer $\boldsymbol{x}^{\star}$. If $T^{k}(\{\boldsymbol{x}^{\star}\})$ is a null set for every $k$ then so is $S$ since the countable union of null sets is a null set.

### v2

@@ Appendix D Additional Convergence Result for Generalized Polyak Stepsize
$C^{-1}(y)$ $=\begin{cases}2y,&\text{ if }y\geq(\gamma K_{\nu})^{\frac{1+\nu}{1-\nu}}\\ (2\gamma K_{\nu}y)^{\frac{1+\nu}{2}},&\text{ if }y<(\gamma K_{\nu})^{\frac{1+\nu}{1-\nu}}\end{cases}$
@@ Appendix E Proofs for Section 6
Definition 7 (Lusin $(N^{-1})$ condition).
Let $T:\mathbb{R}^{d}\to\mathbb{R}^{k}$. We define $T^{-1}$ over a set $S\subseteq\mathbb{R}^{k}$ as
$T^{-1}(S)=\{\boldsymbol{x}:T(\boldsymbol{x})\in S\}.$
We say that $T$ satisfies $(N^{-1})$ condition if for every set $E$ of measure zero we have that $T^{-1}(E)$ also has measure zero.
@@
In other words, the set of initializations that can converge to the fixed point has measure zero.
@@
We divide $\mathbb{R}^{n}$ into two sets, $S=\bigcup_{k=1}^{\infty}T^{-k}(\{\boldsymbol{x}^{\star}\})$, and its compliment $^{c}. SS$ represents the points that can exactly reach the unique minimizer $\boldsymbol{x}^{\star}$. If $T^{k}(\{\boldsymbol{x}^{\star}\})$ is a null set for every $k$ then so is $S$ since the countable union of null sets is a null set.